# Jaynes-Cummings model in a unitary fractional-time description

### v1

@@ Abstract
The time-evolution operator derived from the fractional-time Schrödinger equation is considered non-unitary because it fails to preserve the norm of the vector state as time evolves. However, considering the time-dependent non-Hermitian quantum formalism to the time-fractional dynamics, it has been demonstrated that a unitary evolution can be achieved for a traceless two-level Hamiltonian. This is accomplished by considering a dynamical Hilbert space embedding a time-dependent metric operator, with respect to which the system evolves in a unitary manner, allowing for the proper interpretation of standard quantum mechanical probabilities. In this work, we apply the unitary description to the Jaynes-Cummings model in the fractional-time scenario for investigating the modification in terms of the fractional-order parameter $\alpha$ of the well-known dynamical properties, such as the atomic population inversion of the two-level atom, and the atom-field entanglement when the atom starts in its excited state and field is initially in a coherent state.
@@ I Introduction
Many developments at the interface of physics and fractional calculus have been drawing attention. The cornerstone of this approach lies in substituting the $n$-order derivative $\partial_{z}^{n}f(z)$ by fractional-order differential operators denoted as $\mathcal{D}^{\alpha}_{z}f(z)$, which represents fractional derivatives of order $\alpha$ acting on the function $f(z)$ with respect to the variable $z$. The specific definition of the operator depends on the underlying mathematical functions involved [1]. Interesting applications of fractional calculus appears in statistical physics, particularly in the context of continuous-time random walks (CTRW) to model transport phenomena. CTRW offer a versatile framework for describing anomalous diffusion processes, commonly observed in complex systems exhibiting memory effects and non-local interactions. In the spatial domain, fractional derivative leads to the emergence of Lévy flights [2]. These are characterized by jumps with power-law distributed step lengths, resulting in a diffusion process with long-tailed probability densities. Conversely, incorporating a fractional derivative in the temporal domain leads to subdiffusive behavior [3]. This signifies slower-than-expected diffusion, often observed in systems with heterogeneous environments or trapping mechanisms. The choice of which domain (spatial or temporal) to introduce the fractional derivative depends on the specific physical mechanisms governing the transport process.
@@
In quantum mechanics, a branch of quantum physics rooted in fractional calculus, has emerged as a powerful framework for understanding the behavior of quantum systems with non-local, non-Markovian, and long-range interactions. This burgeoning field encompasses diverse areas such as Lévy flights over quantum paths [4], optics [5, 6], $\mathcal{PT}$-symmetric systems [7], the nonlinear variable-order time fractional Schrödinger equation [8], disorder in the vibrational spectra [9], time-dependent quantum potentials [10], and anomalous diffusion in three-level system [11]. An experimental demonstration by Wu et al. [12] investigated spontaneous emission from a two-level atom in anisotropic one-band photonic crystals. They elegantly employed fractional calculus to resolve an unphysical bound state anomaly arising when the resonant atomic frequency deviates from the photonic band gap. This anomaly, characterized by an infinitely long lifetime, vanishes when the emission peak aligns with the band gap [13].
@@
In this realm, the means of characterizing states involves the application of the fractional Schrödinger equation (FSE), which was first introduced by Laskin. Unlike the conventional Schrödinger equation, Laskin’s formulation replaces the standard second-order spatial derivative with a fractional Laplacian operator derived from the Reisz derivative [14, 15, 16, 17]. This adaptation allows the FSE to model non-local interactions and memory effects, which are critical for comprehending the nuanced transport processes within quantum systems. Furthermore, Naber [18] has proposed a fractional-time Schrödinger equation (FTSE) assuming the Caputo fractional derivative in the place of the ordinary time derivative in such a way that the equation is written as
@@
defining the fractional Caputo derivative for $\alpha\in[0,1)$. $\hat{H}_{\alpha}$ represents the fractional Hamiltonian, and $\hbar_{\alpha}$ the fractional-Planck constant used as a scale factor [see [18]], and therefore, we may consider all the variables and Planck constant in Eq. (1) as dimensionless quantities. In Ref. [18], it is argued that the imaginary unit is raised to the same power as the time coordinate by performing a Wick rotation. More details about this issue are discussed in [19]. Solutions for FTSE has been investigated in many settings, including the fractional dynamics of free particles [18], and particles under the influence of $\delta$ potentials [20]. An interesting noteworthy is a mathematical correspondence between the FTSE and the fractional-time diffusion equation [18, 19, 21, 20], viewed as describing a non-Markovian process. This correspondence can be verified by replacing the real-time for the imaginary-time in the fractional diffusion equation [18]. Also, a connection between classical geometric diffusion and quantum dynamics is elucidated in Ref. [22], wherein continuous-time quantum walks are represented as quantum analogs of turbulent diffusion within comb geometry.
@@
Recently, quantum information science has witnessed remarkable growth and development, propelled by advances in both theoretical and experimental capabilities. This promising field has revolutionized our understanding of information processing at the quantum level, such as quantum cryptography [23], quantum teleportation [24], quantum metrology [25], and quantum control [26], or quantum computing [27]. While fractional calculus offers a rich mathematical framework for describing complex phenomena, its application to quantum information problems remains relatively unexplored. In this sense, Zu and coworkers [28, 29] analyze the memory effect role of FTSE in the time-evolution of a single quantum state and quantum entanglement by considering the Jaynes-Cummings (JC) model [30]. The JC model describes the interaction between a single atom and a single light wave trapped in a cavity. This interaction exhibits fascinating quantum phenomena like Rabi oscillations and entanglement, providing insights into fundamental light-matter interactions and laying the groundwork for advancements in quantum technologies. Within the fractional scenario, the two-level system interacting with the light field is investigated in Refs. [31, 32].
@@
In fact, the FTSE generates many undesired results, such as the non-existence of stationary energy levels, non-unitarity of the evolution, and consequently, the non-conservation of probability, as discussed in Ref. [33]. It becomes evident when we transform the FTSE in a usual Schrödinger-like equation with an effective time-dependent non-Hermitian Hamiltonian operator. It can be explicitly verified by applying the Riemann-Liouville derivative operator ${}^{\text{RL}}_{~{}~{}0}\mathcal{D}_{t}^{1-\alpha}$ on both sides of the Eq. (1), and evoking the following property of the fractional differentiation for $\alpha\in(0,1]$ [34]
@@
As mentioned before, the effective Hamiltonian is non-Hermitian, which implies a non-unitary time-evolution of the quantum state. In this sense, different proposals to map the non-unitary fractional evolution operator into a unitary one have been made [35, 33, 36, 37]. Particularly, in Ref. [37], a unitary evolution for a traceless non-Hermitian two-level system evolving under FTSE was established by applying the time-dependent non-Hermitian quantum formalism [38, 39, 40]. Essentially, the map of the non-unitary fractional time-evolution operator into a unitary one by employing non-Hermitian quantum mechanics procedures for time-dependent metrics allows for a proper quantum mechanical interpretation of the fractional dynamics. For more details about time-dependent non-Hermitian systems see Refs. [38, 41, 42, 39, 43, 44, 45, 46].
In this work, we propose a unitary fractional-time evolution of the Jaynes-Cummings (JC) model, guided by the formalism developed in [37]. In Section II, we introduce the JC model and the fractional-time evolution operator by solving the fractional time Schrödinger equation (FTSE). Additionally, we establish a connection with time-dependent non-Hermitian quantum formalism by constructing a time-dependent Dyson map, which relates to the dynamical Hilbert space metric. This approach enables us to derive an equivalent unitary time-evolution operator that describes the system’s dynamics within the conventional quantum mechanical framework. Subsequently, in Sec. III, we analyze the population inversion dynamics of an atom interacting with a coherent field. We observe collapse and revival phenomena for various values of the fractional-order parameter, $\alpha$. Following this, in Section IV, we quantify the entanglement dynamics using the von Neumann entropy, verifying that the system evolves from a separable state to an entangled one. Our conclusions follow in Sec. V.
II Fractional-time dynamics of the Jaynes-Cummings model
II.1 The model
@@
This model serves as a paradigm for comprehending the fundamental processes governing light-matter coupling at the quantum level. The JC model describes the interaction of a two-level atom with a single quantized mode of the radiation field. In an ideal cavity QED experiment, the atom can be viewed as a two-level system ($|g\rangle$ and $|e\rangle$) coupled to a single mode of the field, and the system evolution is determined by the famous JC Hamiltonian of quantum optics. Let us consider the resonant case, where the atomic energy gap between the two-level is equal to the energy radiation field. In the interaction picture, the Hamiltonian of the JC model with the rotating-wave approximation (RWA) is described by
@@
The JC model is specified via the states of both atom and field, where the basis states of the field are the number states $|n\rangle$, with $n=0,1,2,\cdots$. In this case, the bare states $|g,n\rangle$ and $|e,n\rangle$ provide a natural basis for the infinite-dimensional Hilbert space representing the atom-field interaction. The ground state corresponds to the state with the atom in the ground state $|g\rangle$ and no photons in the cavity $|0\rangle$. In this case, we have the relation
@@
which means that spontaneous absorption from the vacuum is forbidden. Furthermore, for each photon number $n$ the bare states pairs, the Hamiltonian couples the states $|e,n\rangle$ and $|g,n+1\rangle$, since
@@
Therefore, the Hamiltonian can be decomposed as a block-diagonal matrix
@@
with $\mu^{(n)}_{\alpha}=\sqrt{n+1}\mu_{\alpha}$, in which represents the Hamiltonian of the system in the two-dimensional subspace $\mathcal{H}_{n}$. Next, we discuss how the system evolves in time under the FTSE.
II.2 Fractional time-evoluiton
@@
that is a non-unitary operator satisfying the initial condition $\hat{U}_{\alpha}(0)=\hat{1}$. In the above equation, the function $E_{\alpha}(x)=\sum_{k=0}^{\infty}x^{k}/\Gamma(\alpha k+1)$ is identified to be the well-known one-parameter Mittag-Leffler function [1].
@@
The nonunitary nature of time evolution when considering the FTSE might be applied to mimic the effects of the environment on quantum systems. However, this approach would be inherently heuristic, as nonunitarity leads to the non-conservation of probability. In order to establish the conventional interpretation of quantum mechanics, our aim is to map the nonunitary fractional time-evolution operator to a unitary one by employing non-Hermitian quantum mechanics techniques with time-dependent metrics. This procedure enables a proper quantum-mechanical interpretation of the fractional time description, utilizing a modified inner product as done in the non-Hermitian framework.
II.3 Unitary fractional-time evolution
Hereafter we apply the results developed in Ref. [37] for the JC model. Remarkably, within the framework of time-dependent non-Hermitian formalism [38, 40], a state undergoing nonunitary evolution, denoted as $|\Psi_{\alpha}(t)\rangle$, can be linked to a state evolving unitarily, $|\psi_{\alpha}(t)\rangle$, via the time-dependent Dyson map, expressed by the following relation:
such a map is assumed to be invertible. In what follows, from the fact that $|\psi_{\alpha}(t)\rangle=\hat{u}_{\alpha}(t)|\psi_{\alpha}(0)\rangle$, the Eq. (13) allows us to obtain the unitary time-evolution operator $\hat{u}_{\alpha}(t)$ in terms of the Dyson map $\hat{\eta}_{\alpha}(t)$ and the non-unitary time-evolution operator $\hat{U}_{\alpha}(t)$ from the equality
Since the nonunitary time-evolution operator $\hat{U}_{\alpha}(t)$ is known, we have to specify the time-dependent Dyson map parameters for mapping the fractional dynamics in a unitary one. It shows that is possible to define a dynamical Hilbert space with a modified inner-product, defined as $\langle\Psi_{\alpha}(t)|\Psi_{\alpha}(t)\rangle_{\Theta_{\alpha}(t)}=\langle\Psi_{\alpha}(t)|\hat{\Theta}_{\alpha}(t)|\Psi_{\alpha}(t)\rangle$, where the fractional-time evolution can be seen as a unitary in according to modified inner product
$=\langle\psi_{\alpha}(0)|\psi_{\alpha}(0)\rangle$
$=\langle\psi_{\alpha}(t)|\psi_{\alpha}(t)\rangle,$
in which $\hat{\Theta}_{\alpha}(t)=\hat{\eta}_{\alpha}^{\dagger}(t)\hat{\eta}_{\alpha}^{\,}(t)$ is the metric operator. This relation reflects that the probability conservation in the fractional-time scenario can be achieved by defining a suitable time-dependent metric with respect to which the state evolves unitarily. Also, this relation means that it is equivalent to mapping the state that evolves nonunitarily in another system that evolves unitarily with respect to the usual metric.
Indeed, the choice of time-dependent Dyson map $\hat{\eta}_{\alpha}(t)$ is not unique, and for this reason, we propose an Hermitian block-diagonal form given by

### v2

# Unitary description of the Jaynes-Cummings model under fractional-time dynamics

@@ Abstract
The time-evolution operator corresponding to the fractional-time Schrödinger equation is nonunitary because it fails to preserve the norm of the vector state in the course of its evolution. However, in the context of the time-dependent non-Hermitian quantum formalism applied to the time-fractional dynamics, it has been demonstrated that a unitary evolution can be achieved for a traceless two-level Hamiltonian. This is accomplished by considering a dynamical Hilbert space embedding a time-dependent metric operator concerning which the system unitarily evolves in time. This allows for a suitable description of a quantum system consistent with the standard quantum mechanical principles. In this work, we investigate the Jaynes-Cummings model in the fractional-time scenario taking into account the fractional-order parameter $\alpha$ and its effect in unitary quantum dynamics. We analyze the well-known dynamical properties, such as the atomic population inversion and the atom-field entanglement, when the atom starts in its excited state and the field in a coherent state.
@@ I Introduction
Many developments at the interface of physics and fractional calculus have been drawing attention. The cornerstone of this approach lies in substituting the $n$-order derivative $\partial_{z}^{n}f(z)$ by fractional-order differential operators denoted as $\mathcal{D}^{\alpha}_{z}f(z)$, which represents fractional derivatives of order $\alpha$ acting on the function $f(z)$ with respect to the variable $z$. The specific definition of the operator depends on the underlying mathematical functions involved [1]. Interesting fractional calculus applications appear in statistical physics, particularly in the context of continuous-time random walks (CTRW) to model transport phenomena. CTRW offers a versatile framework for modeling anomalous diffusion processes, often observed in complex systems with memory effects and non-local interactions. In the spatial domain, the fractional derivative leads to the emergence of Lévy flights [2]. These are characterized by jumps with power-law distributed step lengths, resulting in a diffusion process with long-tailed probability densities. Conversely, incorporating a fractional derivative in the temporal domain leads to subdiffusive behavior [3]. This signifies slower-than-expected diffusion, often observed in systems with heterogeneous environments or trapping mechanisms. The choice of which domain (spatial or temporal) to introduce the fractional derivative depends on the specific physical mechanisms governing the transport process.
@@
A branch of quantum physics rooted in fractional calculus has emerged as a powerful framework for understanding the behavior of quantum systems with non-local, non-Markovian, and long-range interactions. This burgeoning field encompasses diverse areas such as Lévy flights over quantum paths [4], optics [5, 6], $\mathcal{PT}$-symmetric systems [7], the nonlinear variable-order time fractional Schrödinger equation [8], disorder in the vibrational spectra [9], time-dependent (TD) quantum potentials [10], and anomalous diffusion in a three-level system [11]. An experimental demonstration by Wu et al. [12] investigated spontaneous emission from a two-level atom in anisotropic one-band photonic crystals. They elegantly employed fractional calculus to resolve an unphysical bound state anomaly arising when the resonant atomic frequency deviates from the photonic band gap. This anomaly, characterized by an infinitely long lifetime, vanishes when the emission peak aligns with the band gap [13].
@@
In this realm, states are described by the fractional Schrödinger equation (FSE), originally proposed by Laskin. Unlike the conventional Schrödinger equation, Laskin’s formulation replaces the standard second-order spatial derivative with a fractional Laplacian operator, which is based on the Riesz derivative [14, 15, 16, 17]. This modification enables the FSE to model non-local interactions and memory effects, essential for understanding quantum systems’ transport phenomena. Furthermore, Naber [18] has proposed a fractional-time Schrödinger equation (FTSE) assuming the Caputo fractional derivative in the place of the ordinary time derivative in such a way that the equation is written as
@@
defining the fractional Caputo derivative for $\alpha\in[0,1)$. $\hat{H}_{\alpha}$ represents the fractional Hamiltonian, and $\hbar_{\alpha}$ the fractional-Planck constant used as a scale factor (see [18]), and therefore, we may consider all the variables and Planck constant in Eq. (1) as dimensionless quantities. In Ref. [18], it is argued that the imaginary unit is raised to the same power as the time coordinate by performing a Wick rotation. More details about this issue are discussed in [19]. FTSE solutions have been investigated in many settings, including the fractional dynamics of free particles [18] and particles under the influence of $\delta$ potentials [20]. A mathematical correspondence between the FTSE and the fractional-time diffusion equation [18, 19, 20], viewed as describing a non-Markovian process. Also, a connection between classical geometric diffusion and quantum dynamics is elucidated in Ref. [21], wherein continuous-time quantum walks are represented as quantum analogs of turbulent diffusion within comb geometry.
@@
Recent advances in quantum information science, driven by theoretical and experimental breakthroughs, have significantly expanded our understanding of information processing at the quantum level, leading to quantum cryptography [22], quantum teleportation [23], quantum metrology [24], quantum control [25], and quantum computing [26]. While fractional calculus offers a rich mathematical framework for describing complex phenomena, its application to quantum information problems remains relatively unexplored. In this sense, Zu and coworkers [27, 28] analyze the role of the memory effect of FTSE in the time evolution of a single quantum state and quantum entanglement by considering the Jaynes-Cummings (JC) model [29], describing the interaction between a single atom and a single mode cavity. This interaction exhibits fascinating quantum phenomena like Rabi oscillations and entanglement, providing insights into fundamental light-matter interactions and laying the groundwork for advancements in quantum technologies. The two-level system interacting with the light field in the fractional scenario is also investigated in Refs. [30, 31].
Our work proposes a unitary associated with the fractional-time evolution of the JC model, guided by the formalism developed in [32]. The key concepts underlying such formalism are revisited in Section II, where we briefly discuss the nonunitary nature of fractional time-evolution operator stemming from the Caputo derivative and establish a connection with TD non-Hermitian quantum formalism [33, 34]. This connection is created by constructing a TD Dyson map, which is related to the dynamical Hilbert space metric. In Section III, we introduce the JC model and the fractional-time evolution operator by solving the FTSE and obtain an equivalent unitary time-evolution operator that describes the dynamics within the conventional quantum mechanical formulation. Subsequently, in Sections IV and V, we respectively analyze the collapse and revival phenomena, and quantify the entanglement dynamics of an atom interacting with a coherent field through the von Neumann entropy. Our conclusions follow in Section VI.
II Non-hermiticity and the fractional-time scenario
@@
The FTSE may generate many undesired results, such as the non-existence of stationary energy levels, non-unitarity of the evolution, and consequently, the non-conservation of probability, as discussed in Ref. [35]. It becomes evident when we transform the FTSE in a usual Schrödinger-like equation with an effective TD non-Hermitian Hamiltonian operator. It can be explicitly shown by applying the Riemann-Liouville derivative operator ${}^{\text{RL}}_{~{}~{}0}\mathcal{D}_{t}^{1-\alpha}$ on both sides of the Eq. (1), and utilizing the fractional differentiation property for $\alpha\in(0,1]$ [36], expressed as:
@@
Note that the effective Hamiltonian is non-Hermitian, which implies a nonunitary time evolution of the quantum state. In this sense, different proposals have been made to map the fractional nonunitary evolution operator into a unitary one [37, 35, 38, 32]. In particular, in Ref. [32], a unitary evolution for a traceless non-Hermitian two-level system evolving under FTSE was established by applying the TD non-Hermitian quantum formalism [33, 39, 34].
In the framework established in [32], a state undergoing nonunitary evolution, denoted by $|\Psi_{\alpha}(t)\rangle$, may be mapped to a state $|\psi_{\alpha}(t)\rangle$ evolving unitarily via the TD Dyson map $\hat{\eta}_{\alpha}(t)$. This transformation is expressed as follows
such a map is assumed to be invertible. In what follows, from the fact that $|\psi_{\alpha}(t)\rangle=\hat{u}_{\alpha}(t)|\psi_{\alpha}(0)\rangle$ and $|\Psi_{\alpha}(t)\rangle=\hat{U}_{\alpha}(t)|\Psi_{\alpha}(0)\rangle$, the Eq. (6) allows us to obtain the unitary time-evolution operator $\hat{u}_{\alpha}(t)$ in terms of the TD Dyson map and the nonunitary time-evolution operator as being
Assuming the nonunitary time-evolution operator $\hat{U}_{\alpha}(t)$ is known, our primary goal is to identify the TD Dyson map parameters that enable the mapping of nonunitary dynamics to a unitary one. This can be accomplished by consider that $\hat{u}_{\alpha}(t)$ satisfies a Schrödinger-like equation
$i\hbar\partial_{t}\hat{u}_{\alpha}(t)=\hat{h}_{\alpha}(t)\hat{u}_{\alpha}(t),$ (8)
and employing Eqs. (5), (7), and (8) to obtain the Hermitian Hamiltonian $\hat{h}_{\alpha}(t)$, which is given by
$\hat{h}_{\alpha}(t)=\hat{\eta}_{\alpha}(t)\hat{H}_{\alpha}^{\text{eff}}(t)\hat{\eta}_{\alpha}^{-1}(t)+i\hbar\partial_{t}\hat{\eta}_{\alpha}(t)\hat{\eta}_{\alpha}^{-1}(t),$ (9)
being $\hat{H}_{\alpha}^{\text{eff}}(t)$ the effective Hamiltonian into Eq. (5):
$\hat{H}_{\alpha}^{\text{eff}}(t)=\frac{i^{1-\alpha}\hbar}{\hbar_{\alpha}}\hat{H}_{\alpha}{\,}^{\text{RL}}_{~{}~{}0}\mathcal{D}_{t}^{1-\alpha}\hat{U}_{\alpha}(t).$ (10)
Note that the effective Hamiltonian is not a pseudo-Hermitian operator and, therefore, it cannot be an observable [33]. However, if a TD Dyson map can be found that transforms the non-Hermitian Hamiltonian into a Hermitian one, it is possible to construct a dynamical Hilbert space equipped with a modified inner product, defined as $\langle\Psi_{\alpha}(t)|\Psi_{\alpha}(t)\rangle_{\Theta_{\alpha}(t)}=\langle\Psi_{\alpha}(t)|\hat{\Theta}_{\alpha}(t)|\Psi_{\alpha}(t)\rangle$, where the fractional-time evolution can be seen as a unitary in according to modified inner product
$=\langle\psi_{\alpha}(0)|\psi_{\alpha}(0)\rangle$
$=\langle\psi_{\alpha}(t)|\psi_{\alpha}(t)\rangle,$
in which $\hat{\Theta}_{\alpha}(t)=\hat{\eta}_{\alpha}^{\dagger}(t)\hat{\eta}_{\alpha}^{\,}(t)$ is the positive metric operator. The above relation reflects that the probability conservation in the fractional-time scenario can be achieved by defining a suitable TD metric with respect to which the state evolves unitarily. Moreover, this relation means that it is equivalent to mapping the state that evolves nonunitarily in another system that evolves unitarily to the usual metric. For more details about TD non-Hermitian systems see Refs. [33, 40, 41, 34, 39, 42, 43, 44, 45].
III Fractional-time dynamics of the Jaynes-Cummings model
III.1 The model
@@
Now we analyze the JC model in the fractional-time scenario revised in the previous section. The JC model is a paradigmatic framework for understanding quantum light-matter interactions [29]. It describes the interaction of a two-level atom with a single quantized mode of the radiation field. In an ideal cavity QED experiment, the atom can be viewed as a two-level system ($|g\rangle$ and $|e\rangle$) interacting with a single mode of the field. This interaction is described by the JC Hamiltonian. Consider the resonant case, in which the energy difference between the two atomic levels exactly matches the energy of the radiation field. Within the rotating-wave approximation (RWA), the JC Hamiltonian in the interaction picture is given by
@@
The model is specified via the states of both atom and field, where the basis states of the field are the number states $|n\rangle$, with $n=0,1,2,\cdots$. In this case, the bare states $|g,n\rangle$ and $|e,n\rangle$ provide a natural basis for the infinite-dimensional Hilbert space representing the atom-field interaction. The ground state corresponds to the state with the atom in the ground state $|g\rangle$ and no photons in the cavity $|0\rangle$. In this case, we have the relation
@@
which means that spontaneous absorption from the vacuum is forbidden. Furthermore, for each photon number $n$, the Hamiltonian couples the bare states pairs $|e,n\rangle$ and $|g,n+1\rangle$, since
@@
The Hamiltonian can be decomposed as a block-diagonal matrix
@@
which represents the Hamiltonian of the system in the two-dimensional subspace $\mathcal{H}_{n}$, with $\mu^{(n)}_{\alpha}=\sqrt{n+1}\mu_{\alpha}$. A concise treatment of the JC model can be found in Ref. [46]. Next, we discuss how the system evolves under the FTSE and can be described into the unitary framework.
III.2 Fractional time-evolution
@@
that is a nonunitary operator satisfying the initial condition $\hat{U}_{\alpha}(0)=\hat{1}$. In the above equation, the function $E_{\alpha}(x)=\sum_{k=0}^{\infty}x^{k}/\Gamma(\alpha k+1)$ is identified to be the well-known one-parameter Mittag-Leffler function [1].
@@
While Naber’s approach to the FTSE is often regarded as non-physical, it remains a widely used framework for mathematical investigations. However, in the specific case of $\alpha=1/2$, the solution described by the Mittag-Leffler function acquires a meaningful physical interpretation, as discussed in Ref. [47]. This unique feature –where the $\alpha=1/2$ fractional order gains physical significance– also arises in other approaches, as proposed in Refs. [48, 49], where the fractional derivative emerges naturally within their respective frameworks.
The nonunitary nature of time evolution when considering the FTSE might be applied to mimic the effects of the environment on quantum systems. However, this treatment would be inherently heuristic, as nonunitarity leads to the non-conservation of probability. For a treatment of open quantum fractional dynamics, we suggest Ref. [50], where a framework describing the nonunitary evolution of the density operator that is trace-preserving and completely positive for any initial condition is developed. Nevertheless, in this work, we are interested in establishing the conventional interpretation of quantum mechanics, by mapping the nonunitary fractional time-evolution operator to a unitary one by employing non-Hermitian quantum mechanics techniques with TD metrics [32]. This procedure enables a proper quantum-mechanical interpretation of the fractional-time description, utilizing a modified inner product as done in the non-Hermitian framework [33, 34].
III.3 Unitary evolution
Indeed, the choice of TD Dyson map is not unique. To address this, we propose a Hermitian time-dependent Dyson map, $\hat{\eta}_{\alpha}(t)$, in block-diagonal form, expressed as: